A firm has a total production cost of $200,000. Its average fixed cost is $120 and its average variable


cost is $80. Calculate the firm’s total fixed costs.

1
Expert's answer
2021-12-07T10:24:39-0500

Solution

Total Production Cost "=TC=" $"200,000"

Average Fixed Cost "=AFC=" $"120"

Average Variable Cost "=AVC=" $"80"

Quantity "=q"


"TC=(AFC+AVC)*q"

"200,000=(120+80)*q"

"200,000=200q"

"q=" 1,000 units


The firm’s total fixed costs(TFC):

"TFC=AFC*q"

"=" $"120*1,000"

"=" $"120,000"

Thus, the firm's total fixed costs is $120,000